OpenStudy (anonymous):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. Vertex at (-5,0) and co-vertex at (0,4) A. x^2/25 + y^2/16 =1 B. x^2/16 + y^2/25 =1 C. x^2/5 + y^2/4 =1 D. x^2/4 + y^2/5 =1

OpenStudy (amistre64):

might be simpler to just square the x and y parts and match them with the options

OpenStudy (amistre64):

when x=-5 and y=0; x^2/ ? = 1?

OpenStudy (mertsj):

The center is at the origin. The vertex is (-5,0) so a=5. The covertex is at (0,4) of b = 4 Use those facts and fill in this equation: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]